OpenStudy (anonymous):

Given polygon BRAIN~POWEQ what is the perimeter of BRAIN? A. 23.50 B. 58.75 C. 35.25 D. 57.25

jimthompson5910 (jim_thompson5910):

first you need to know the value of y

jimthompson5910 (jim_thompson5910):

how do you find this

OpenStudy (anonymous):

using slope intercept right

jimthompson5910 (jim_thompson5910):

no you just set up a proportion

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

(y+5)/4.5 = 7.5/3 y = ??

OpenStudy (anonymous):

I got y=6.25

jimthompson5910 (jim_thompson5910):

me too

jimthompson5910 (jim_thompson5910):

so if y = 6.25, then y+5 = 6.25+5 = 11.25

jimthompson5910 (jim_thompson5910):

so the length of BR goes from y+5 to 11.25

jimthompson5910 (jim_thompson5910):

you now have all the info you need to find the perimeter of BRAIN

OpenStudy (anonymous):

so what answer would it Be?

jimthompson5910 (jim_thompson5910):

you tell me

jimthompson5910 (jim_thompson5910):

what did you get

jimthompson5910 (jim_thompson5910):

you just add up all the sides in the figure

OpenStudy (anonymous):

I got A.

jimthompson5910 (jim_thompson5910):

the five sides are: 10,11.25,7.5,12.5,17.5

OpenStudy (anonymous):

ohh 58.75

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

I got one more

OpenStudy (anonymous):

find the value of Y in the similar figures (round to the nearest tenth) A. 12 B. 6 C. 13.3 D. 7.3 One sec gota draw it quick

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

y/20 = 4/6 y = ??

OpenStudy (anonymous):

I got y=40/3

OpenStudy (anonymous):

then simplify that and get 13.3

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

yay!!! thanks for working with me tonight!

jimthompson5910 (jim_thompson5910):

you're welcome